OpenStudy (anonymous):

what is the surface area of the cone? with a radius of 7cm and a height of 24cm. a-175 pi cm^2 b-224 pi cm^2 c-273 pi cm^2 d- 147 pi cm^2

OpenStudy (sleepyhead314):

Hello and Welcome to OpenStudy! :) the surface area of a cone = pi(radius)(√(height^2 + radius^2)) + pi(radius)^2 so plug in radius = 7 and height = 24

OpenStudy (anonymous):

i dont understand i did that not giving the answer i need

OpenStudy (sleepyhead314):

sorry if it was a little confusing :/ try 7*(√(24^4+7^2))pi + (7^2)pi 7*(25)pi + 49pi = ?

OpenStudy (anonymous):

703.71675

OpenStudy (sleepyhead314):

Correct :) That's what I got ^_^

OpenStudy (anonymous):

but that is not an option for my answer ugh?

OpenStudy (sleepyhead314):

hmmm what are your options then?

OpenStudy (anonymous):

a= 175 pi cm^2 b= 224 pi cm^2 c=273 pi cm^2 d= 147 pi cm^2

OpenStudy (sleepyhead314):

ah right, so they want you to keep pi in there so you can either divide the 703.7167544 thing by pi or (7*25) + (49) = ?